# Fairness and Bias in Multimodal : A Survey

### v1

@@ Abstract
The importance of addressing fairness and bias in artificial intelligence (AI)systems cannot be over-emphasized. Mainstream media has been awashed with news of incidents around stereotypes and bias in many of these systems in recent years. In this survey, we fill a gap with regards to the minimal study of fairness and bias in Large Multimodal Models (LMM s) compared to Large Language Models (LLM s), providing 50 examples of datasets and models along with the challenges affecting them; we identify a new category of quantifying bias (preuse), in addition to the two well-known ones in the literature: intrinsic and extrinsic; we critically discuss the various ways researchers are addressing these challenges. Our method involved two slightly different search queries on Google Scholar, which revealed that 33,400 and 538,000 links are the results for the terms "Fairness and bias in Large Multimodal Models" and "Fairness and bias in Large Language Models", respectively. We believe this work contributes to filling this gap and providing insight to researchers and other stakeholders on ways to address the challenge of fairness and bias in multimodal AI.
@@ 1 Introduction
Fairness and bias are very important topics that cut across many domains in the society. The rapid advancements in the research and applications of artificial intelligence (AI)have made them even more compelling in recent times, such that many studies have emerged on them Frankel and Vendrow (2020); Booth et al. (2021); Adewumi et al. (2022); Teo et al. (2024). One important gap in the literature, however, is that there is minimal study or survey on "Fairness and bias in Large Multimodal Models." By multimodal AI, we mean the dataset or AI model that can take one or more modalities as input and/or another as output. This gap is evidenced by the fact that there is a dearth of work on the topic. For example, an exact-phrase search on Google Scholar returns 33,400 links compared to 538,000 links for "Fairness and bias in Large Language Models." [^1] June 10, 2024 This implies more than 16 times the result for the former. Filtering the publication year range to 2014-2024 reduces the links to 17,700 and 20,300, respectively. We intend to contribute in filling that gap in this work.
@@
The two terms, fairness and bias, are strongly related but fairness is concerned with equality and justice while bias is concerned with systematic error, which may arise from human prejudices Booth et al. (2021); Alkhaled et al. (2023). For the purpose of this survey, fairness may be defined as equal representation with regards to a given Sensitive Attribute (SA)Hutchinson and Mitchell (2019); Frankel and Vendrow (2020). Hence, we may consider a generative artifical intelligence (GenAI)to be fair if it generates both male and female samples with equal probabilities, with regards to the SA gender Teo et al. (2024). Bias is a (non-random) systematic error in a measurement resulting in a difference in accuracy in one group compared to another, given the ground truth Booth et al. (2021); Scheuneman (1979). We acknowledge there are other quantitative definitions of fairness and bias, as noted by Hutchinson and Mitchell (2019) and Weidinger et al. (2022).
@@
It appears the emergence of big data, which has brought rapid advancement in the state-of-the-art (SotA), also brought along the increase in poor quality content and prediction, such as the increased criminal prediction for Black and Latino people observed by Birhane et al. (2024a). Similar issues are observed across many domains, including healthcare, employment, forensics, criminal justice, credit scoring, and computational social science, among others Liang et al. (2021); Ferrara (2023); Landers and Behrend (2023); Han (2023). According to Wolfe et al. (2023), the model VQGAN-CLIP, similarly to Stable Diffusion, generated sexualized images for the harmless prompt "a 17 year old girl," 73% of the time. The comparison to a similar prompt with the term "girl" replaced with "boy" shows a sharp contrast.
@@
    We fill the gap with a comprehensive survey of fairness and bias across a wide spectrum of LMM s, LLM s, and multimodal datasets, providing 50 examples of datasets and models in a structured way, along with the challenges affecting them.
@@
    We identify a new category of quantifying bias (i.e. preuse), in addition to the two well-known ones in the literature: intrinsic and extrinsic Ramesh et al. (2023); Cabello et al. (2023).
@@
The rest of this paper is organised as follows. In the following Section (2), we discuss fairness and bias in AI, highlighting some of the theories. In Section 3, we explain the method used for this survey. In Section 4, we focus on various works along the two paradigms of LMM s and LLM s, discussing the datasets and models in the literature. In Section 5, we discuss widely on the methods to evaluate fairness and bias, the datasets for evaluation, and the debiasing strategies. We conclude the survey in Section 6 with a summary and possible future work.
@@ 2.1 Concepts of Fairness and Bias
A discussion about every possible type of fairness or bias is beyond the scope of this study, however, we refer readers to Mehrabi et al. (2021) for some of the discussion. In this section, we highlight some of the concepts around fairness and bias found in the Social Science literature.
@@ 2.2 Consequences & Legal Perspectives
Fredrickson and Roberts (1997) confirms that objectification victimises the subject and may result in habitual body monitoring, thereby increasing mental health risks, sexual dysfunction, eating disorders, and depression. This unhealthy reality is also confirmed by Swim et al. (2001). They realized that sexist incidents occur more against women and have negative emotional consequences for them. Some of these incidents are traditional gender role stereotypes, degrading remarks, and sexual objectification, which are found in the data used for training AI models. For details about the mechanics of training language models, we refer readers to Radford et al. (2019) and Hoffmann et al. (2022). It is not surprising, therefore, that the use of these models cause the same negative effects for those affected. Hence, fairness and bias are not only ethical or moral issues but have legal implications Landers and Behrend (2023). In the United States (US), disparate treatment because of sensitive attributes is unlawful Berry (2015); Hutchinson and Mitchell (2019); Meng et al. (2022). This is also the case in some other countries Zafar et al. (2017).
3 Method
To address the gap identified earlier, our method involved the simple approach of filtered search on Google scholar with two slightly differently-worded phrases: (1) "Fairness and Bias in Large Multimodal Models" and (2) "Fairness and Bias in Large Language Models" and they returned 17,700 and 20,300 links,1 respectively, over many pages. The search was filtered to the period 2014-2024 during which deep learning made significant progress. Apparently, there are fewer scientific papers on the former. We compare with "Language Models" only because these, with the introduction of the Transformer architecture Vaswani et al. (2017), have influenced computer vision Yuan et al. (2021) and serve as core component for many multimodal AI.
We captured archived papers (e.g. arXiv) but then thoroughly reviewed pair-reviewed papers that were returned on the first 10 pages (100 links) for each search text and identified the publishers (or venues), as listed in Table 1. We note that the more relevant papers turn up on the first page of the search while the quality of results degrade as one progresses through the pages. From the Table, it can be observed that there are 36 pair-reviewed papers on multimodal compared to 39 on language, among the 200 papers. We note that Google Scholar sometimes returns the arXiv version of a paper instead of the peer-reviewed version, hence, it’s very useful to include the data from arXiv. In total, 55 papers were published under multimodal compared to 88 for language. Each paper was reviewed for their contributions, including the datasets, models, possible solutions proffered on fairness and bias, and other relevant discussions.
Table 1: Distribution of scientific papers (2014-2024)
@@ 3 Method
1 IEEE 11 2
2 ELSEVIER 2 2
3 ACM 9 9
4 Springer 1 1
5 NeurIPS 4 8
6 Nature 3 1
7 MDPI - 1
8 MLR 3 5
9 PKP 2 -
@@ 15 JMLR - 2
16 Science - 1
Sub-total 36 39
17 arXiv 19 49
Total 55 88
4 Fairness and Bias in LLM s and LMM s
In this section we review some LMM s and LLM s and the fairness and bias challenges they have. Tables 2 and 3 summarize some relevant datasets and the models, respectively. All the 25 datasets identified have their challenges and by extension the 25 AI models they are trained on. Some of these challenges include stereotypes, porn, misogyny, racial, gender, religious, cultural, age, and demographic biases. It is commonly agreed AI models learn much of their bias from the data they are trained on and many datasets, especially those for pretraining, are from the Internet, which contains a diverse spectrum of content Wolfe and Caliskan (2022).
@@ 4.1 LMM s
Table 2: Summary of Some Datasets and Their Fairness & Bias Challenges (Data in the second part of the table are usually used in downstream tasks).
# Dataset Modality Some Challenges
@@
10 The Pile Gao et al. (2020) Text Religious, racial, gender biases Gao et al. (2020)
@@
22 Multi30K Elliott et al. (2016) Text & Vision Racial bias Wang et al. (2022)
23 MIMIC-IV Johnson et al. (2023) Text & Vision Ethnic, racial, marital status biases Meng et al. (2022)
@@
25 Twitter corpus Huang et al. (2020) Text Age, gender, racial biases Huang et al. (2020)
Wolfe et al. (2023) found evidence of sexual objectification bias in models based on Contrastive Language-Image Pretraining (CLIP). The 9 CLIP models that were investigated were trained on internet-wide web crawls. CLIP is known to be quite accurate Radford et al. (2021), however, it also appears to have scaled the biases inherent in its training data. Also, Wolfe and Caliskan (2022) found that more than Latino, Asian or Black, White persons are more associated with collective in-group words in embeddings from CLIP Radford et al. (2021), SLIP Mu et al. (2022), and BLIP Li et al. (2022), as measured with EATs. For a definitive assessment, their work would have benefited from additional experiments involving data of people outside the United States (US), since they used Chicago Face Database (CFD)Ma et al. (2015). CFD is a dataset of 597 people recruited in the U.S. Similarly, Teo et al. (2024) found that Stable Diffusion exhibits gender bias on slight changes to the prompts.
Besides work on text and text-visual multimodal systems or data, there are some work on audio-visual systems or data Fenu and Marras (2022). In the work by Fenu and Marras (2022), they perform comparative analysis on audio-visual speaker recognition systems, using fusion at the model step. They found that the highest accuracy and the lowest disparity across groups are achieved compared to unimodal systems.
In other works, Peña et al. (2023) evaluated AI-based recruitment for multimodal data but in a fictitious case study, which may limit its generalizability in real-world applications. Booth et al. (2021) performed a case study of automated video interviews and found that combining more than one modality increases bias and reduces fairness, similarly to what happens when scaling crawled data for training models. Cabello et al. (2023) investigated fairness and bias in LMM s. Other researchers investigated the impact of multimodal data/models on personlity assessment Yan et al. (2020), cyberbullying Alasadi et al. (2020), health records Meng et al. (2022) and more Birhane et al. (2021); Ferrara (2023); Zhang et al. (2022)
4.2 LLM s
@@
# LMM s Modality Training Data Some Challenges
1 VQGAN-CLIP Crowson et al. (2022) Text & Vision WIT & ImageNet Misogyny, racial bias Pagliai et al. (2024); Wolfe and Caliskan (2022)
2 DALL-E 2 Ramesh et al. (2021) Text & Vision Conceptual Captions Occupational stereotypes, gender bias Ramesh et al. (2022); Mandal et al. (2023)
@@
5 SLIP Mu et al. (2022) Text & Vision YFCC100M Racial bias Wolfe and Caliskan (2022)
6 CLIP Radford et al. (2021) Text & Vision WIT Racial bias Radford et al. (2021); Wolfe and Caliskan (2022)
7 BLIP Li et al. (2022) Text & Vision MS COCO, Conceptual Captions & LAION-400M Racial, gender & age biases Wolfe and Caliskan (2022); Ruggeri and Nozza (2023)
8 PaliGemma Text & Vision WebLI, CC3M-35L , WIT Porn, offensive language222ai.google.dev/gemma/docs/paligemma/model-card
@@
11 BEiT Wang et al. (2023) Text & Vision Conceptual 12M, ImageNet-21K, Wikipedia Gender, cultural biases Brinkmann et al. (2023)
@@
23 Aurora-M Nakamura et al. (2024) Text The Pile Offensive language, religious, racial, gender biases Gao et al. (2020); Nakamura et al. (2024)
24 Mixtral-8x7B Jiang et al. (2024) Text web text Stereotypes, racial, gender, occupational biases Jiang et al. (2024); Aich et al. (2024)

### v2

@@ Abstract
The importance of addressing fairness and bias in artificial intelligence (AI)systems cannot be over-emphasized. Mainstream media has been awashed with news of incidents around stereotypes and other types of bias in many of these systems in recent years. In this survey, we fill a gap with regards to the relatively minimal study of fairness and bias in Large Multimodal Models (LMM s) compared to Large Language Models (LLM s), providing 50 examples of datasets and models related to both types of AI along with the challenges of bias affecting them. We discuss the less-mentioned category of mitigating bias, preprocessing (with particular attention on the first part of it, which we call preuse). The method is less-mentioned compared to the two well-known ones in the literature: intrinsic and extrinsic mitigation methods. We critically discuss the various ways researchers are addressing these challenges. Our method involved two slightly different search queries on two reputable search engines, Google Scholar and Web of Science (WoS), which revealed that for the queries ‘Fairness and bias in Large Multimodal Models’ and ‘Fairness and bias in Large Language Models’, 33,400 and 538,000 links are the initial results, respectively, for Scholar while 4 and 50 links are the initial results, respectively, for WoS. For reproducibility and verification, we provide links to the search results and the citations to all the final reviewed papers. We believe this work contributes to filling this gap and providing insight to researchers and other stakeholders on ways to address the challenges of fairness and bias in multimodal and language AI.
@@ 1 Introduction
Fairness and bias are very important topics that cut across many domains in the society. The rapid advancements in the research and applications of artificial intelligence (AI)have made them even more compelling in recent times, such that many studies have emerged on them Frankel and Vendrow (2020); Booth et al. (2021); Adewumi et al. (2022); Teo et al. (2024). One important gap in the literature, however, is that there is relatively minimal study or survey on ‘Fairness and bias in Large Multimodal Models.’ By multimodal AI, we mean the datasets or AI models that can take one or more modalities as input and/or another as output. This gap is evidenced by the fact that there are fewer works around the topic. For example, a query search on Google Scholar returns 33,400 links compared to 538,000 links for ‘Fairness and bias in Large Language Models’ (where the first query search is equivalent to the boolean operation fairness AND and AND bias AND in AND large AND multimodal AND models). [^1] July 10, 2024 This implies more than 16 times the result compared to the former. However, filtering the publication year range to 2014-2024 reduces the links to 17,200 and 19,300, respectively. We intend to contribute in filling that gap in this work.
@@
The two terms, fairness and bias, are strongly related but fairness is concerned with equality and justice while bias is concerned with systematic error, which may arise from human prejudices Booth et al. (2021); Alkhaled et al. (2023). For the purpose of this survey, fairness may be defined as equal representation with regards to a given Sensitive Attribute (SA)Hutchinson and Mitchell (2019); Frankel and Vendrow (2020). Hence, we may consider a generative artifical intelligence (GenAI)to be fair if it generates both male and female samples with equal probabilities, with regards to the SA gender Teo et al. (2024). Bias is a (non-random) systematic error in a measurement resulting in a difference in accuracy in one entity compared to another, given the ground truth Booth et al. (2021); Scheuneman (1979). We acknowledge there are other quantitative definitions of fairness and bias, as noted by Hutchinson and Mitchell (2019) and Weidinger et al. (2022).
@@
It appears the emergence of big data, which has brought rapid advancement in the state-of-the-art (SotA), also brought along the increase in poor quality content and prediction, such as the increased criminal prediction for Black and Latino people observed by Birhane et al. (2024a). Similar issues are observed across many domains, including healthcare, employment, forensics, criminal justice, credit scoring, and computational social science, among others Liang et al. (2021b); Ferrara (2023a); Landers and Behrend (2023); Han (2023). According to Wolfe et al. (2023), the model VQGAN-CLIP, similarly to Stable Diffusion, generated sexualized images for the harmless prompt "a 17 year old girl," 73% of the time. The comparison to a similar prompt with the term "girl" replaced with "boy" shows a sharp contrast.
In related work, many of the relatively recent surveys on fairness and bias in AI appear to have been of a general nature or focused on other areas. Pagano et al. (2023) focused their attention on ML generally and the 5-year period between 2017 and 2022, thereby missing the nuances and some of the details related to natural language processing (NLP)and multimodal AI. Mehrabi et al. (2021) surveyed applications that have exhibited bias in different domains, listed sources of bias in these applications and created a taxonomy for fairness definitions. On the other hand, Le Quy et al. (2022) paid attention to benchmark tabular datasets for fairness, analysing relationships between different protected and class attributes. Balayn et al. (2021) focused on data bias in data engineering and management research, arguing for the enforcement of fairness requirements and constraints on the data for training and evaluating systems. Their survey method is limited in that it restricted part of its literature search between 2019 and 2020. In our work, in addition to discussing datasets that make AI models biased and the datasets for evaluating bias and fairness, we discuss other important related areas, such as the mitigation strategies.
Blodgett et al. (2020) surveyed over 140 articles about bias in NLP and realized that the stated motivations are usually inconsistent and vague, and the articles do not engage with the broader applicable literature that is external to NLP. They, therefore, made 3 recommendations for those in this field: (1) establish the relationships in language and social hierarchies by referring to the broader literature outside of NLP, (2) be explicitly clear on why the system described as biased is harmful, and (3) evaluate the language of people affected by the biased systems. Their survey was restricted to only articles on text-based NLP, thereby excluding speech or multimodal AI, and limited to articles before May 2020. However, we follow their recommendations and our work answers some of the research questions identified in their work. For example, Section 4 and subsection 5.2.1 address the question How are datasets collected?. The survey by Sun et al. (2019) focused on recognizing and mitigating NLP gender bias and the authors discussed this based on four forms of representation bias: denigration, stereotyping, recognition, and under-representation. In Haltaufderheide and Ranisch (2024), they identified ethical issues around fairness and bias with LLM s in medicine and healthcare. Additional works that have surveyed fairness and bias in LLM s include Bender et al. (2021); Meade et al. (2022); Gallegos et al. (2023); Chang et al. (2024); Myers et al. (2024).
@@
    We fill the gap with a comprehensive survey of fairness and bias across a wide spectrum of LMM s, LLM s, and multimodal datasets, providing 50 examples of datasets and models in a structured way, along with the challenges of bias affecting them.
@@
    We discuss the less-mentioned category of mitigating bias in NLP (i.e. preprocessing), though more common in general ML Mehrabi et al. (2021); Pagano et al. (2023), with particular attention on the first part of it, which we call preuse. The other two well-known mitigation methods in the literature are the intrinsic (or in-processing) and extrinsic (or post-processing) methods Ramesh et al. (2023); Cabello et al. (2023).
@@
The rest of this paper is organised as follows. In the following Section (2), we highlight some of the theories around fairness and bias in AI. In Section 3, we explain the method used for this survey. In Section 4, we focus on various works along the two paradigms of LMM s and LLM s, discussing the datasets and models in the literature. In Section 5, we discuss widely on the methods to evaluate fairness and bias, the datasets for evaluation, and the debiasing strategies. We conclude the survey in Section 6 with a summary and possible future work.
@@ 2 Fairness and Bias in AI
2.1 Implicit and Explicit Bias
There is distinction between implicit and explicit bias such that, given two sets of terms that express the bias axis (e.g. gender or race), one set of male gender terms could be $\{S_{1}\}$ = {dad, man} and the other set of female terms be $\{S_{2}\}$ = {mum, woman}. Implicit bias is sufficiently specified with both lists. However, explicit bias requires two additional sets of attributes $\{A_{1}\}$ = {engineer, doctor} and $\{A_{2}\}$ = {caregiver, carer} that express the terms to which the earlier gendered terms exhibit association, albeit to different levels. Hence, a gender biased system could result in male terms in $\{S_{1}\}$ being strongly associated with attributes of career terms $\{A_{1}\}$ compared to female terms $\{S_{2}\}$, which could be strongly associated with attributes of home-related terms $\{A_{2}\}$ Friedrich et al. (2021).
Besides these two broad distinction of bias, there are many types of bias, depending on the philosophical or social perspectives that may be taken. Hence, a discussion about every possible type of fairness or bias is beyond the scope of this study but we refer readers to Mehrabi et al. (2021), Van der Wal et al. (2024) and Navigli et al. (2023) for some of the many types that may be listed.
@@ 2.2 Concepts of Fairness and Bias
In this section, we highlight a few of the concepts around fairness and bias found in the Social Science literature.
@@ 2.3 Consequences of Bias
Fredrickson and Roberts (1997) confirms that objectification victimises the subject and may result in habitual body monitoring, thereby increasing mental health risks, sexual dysfunction, eating disorders, and depression. This unhealthy reality is also confirmed by Swim et al. (2001). They realized that sexist incidents occur more against women and have negative emotional consequences for them. Some of these incidents are traditional gender role stereotypes, degrading remarks, and sexual objectification, which are found in the data used for training AI models. For details about the mechanics of training language models, we refer readers to Radford et al. (2019) and Hoffmann et al. (2022). It is not surprising, therefore, that the use of these models cause the same negative effects for those affected. Hence, fairness and bias are not only ethical or moral issues but have legal implications Landers and Behrend (2023). In the United States (US), disparate treatment because of sensitive attributes is unlawful Berry (2015); Hutchinson and Mitchell (2019); Meng et al. (2022a). This is also the case in some other countries Zafar et al. (2017).
3 Methodology
In order to have a fair and thorough survey for the stated objective in Section 1, we followed the general guidelines recommended for conducting a systematic literature review, which is founded on a rigorous and auditable methodology Kitchenham (2004); Brereton et al. (2007). We used two common scientific search engines: Google Scholar and Web of Science (WoS). Both are advantageous because they index the main literature databases or publishers, including the Institute of Electrical and Electronics Engineers (IEEE), Association for Computing Machinery (ACM), Conference on Neural Information Processing Systems (NeurIPS), Multidisciplinary Digital Publishing Institute (MDPI), Public Knowledge Project (PKP), Massachusetts Institute of Technology Press (MIT Press), Public Library of Science (PLoS), Association of Computational Linguistics (ACL), AI Access Foundation (AIAF), Higher Education Press (HE Press), Association of Measurement and Evaluation in Education and Psychology (EPODDER), New York University School of Law (NYUSL), Academic Conferences and Publishing International (ACPI), Machine Learning Research Press (MLRP), Journal of Machine Learning Research (JMLR), and Society of Photographic Instrumentation Engineers (SPIE).
WoS provides a second assessment to Scholar because of the limitations of the latter, including its inclusion of non-peer-reviewed links or possible predatory journals in its results of primary studies (or papers) Foltỳnek et al. (2019). The two search engines are identified as largely reliable, helping to mitigate the risk of incompleteness in the search Foltỳnek et al. (2019). For both, we employed a similar multi-phase search process itemized below, which helped to bring refinement to each search and the final results. We are confident this approach returned the most applicable results for the purpose of this study.
  1. 1.
    In the first phase, we designed the following two search queries based on our objective.
    • Fairness and Bias in Large Multimodal Models
    • Fairness and Bias in Large Language Models
    The boolean operation, in both search engines, for the first is equivalent to fairness AND and AND bias AND in AND large AND multimodal AND models while the second is fairness AND and AND bias AND in AND large AND language AND models.
  2. 2.
    In the second phase, we filtered the results based on the time period of just over 10 years (2014 to 2024) and relevance (after a review of paper titles and abstracts for over 200 papers). For example, one result involved an article about ‘executive coaching programs’ that has nothing to do with AI or social bias. All the filtered papers are in English. In addition, we corrected for misplaced results (i.e. papers that turned up in one search result though they belong in the other; about 15 in multimodal belonging in language and 9 in language belonging in multimodal, for Scholar). We removed papers published in journals in Beall’s List of Potential Predatory Journals and Publishers,222beallslist.net/standalone-journals if present.
  3. 3.
    In the third phase, we critically reviewed the papers for their contributions, including the datasets, models, possible solutions proffered on fairness and bias, and other relevant discussions.
More concretely, for WoS, we searched ‘All Fields’ of the documents ‘Core Collection’ across all ‘Editions’ without initially restricting the year of publication. For Scholar, the first phase returned 33,400 and 538,000 result links over many pages for the first and second queries, respectively, while returning 4 and 50 for WoS.1 Filtering Scholar, based on the time period, returned 17,200 and 19,300 links, which finally reduced to 69 and 101 at the end of the second phase for the first and second terms, respectively, while returning 8 and 44 for WoS. It is noteworthy that for Scholar, there were equally 100 links to start with for both queries. The boolean search operation involved all the individual words in any arbitrary order without case sensitivity and narrowed the search to mostly relevant documents. We compare ‘Multimodal Models’ with only ‘Language Models’ in the search terms because the latter, with the introduction of the Transformer architecture Vaswani et al. (2017), have influenced computer vision Yuan et al. (2021) and they serve as important components for many multimodal AI. Furthermore, we recognize that multimodal implies the combination of ‘language’ (or ‘speech’) and ‘vision’ but we did not use this distinction in our multimodal search because (1) the boolean equivalent can result in false positives, (2) even when we attempted it, we had less than 50% links in result compared to the second term of only ‘language’ search, and (3) ‘multimodal’ is a standard term in the field.
Table 1: Distribution of scientific papers on Google Scholar. (Filtered total is limited to the year range 2014 - 2024 and to the first 100 links per search query, excluding irrelevant links.)
@@ 3 Methodology
Unfiltered total 33,400 538,000
1 IEEE 10 -
2 Elsevier 3 3
3 ACM 14 13
4 Springer 7 5
5 NeurIPS 4 11
6 Nature 3 5
7 MDPI 2 1
8 MLRP 2 4
9 PKP 1 1
@@ 14 JMLR - 2
15 ACL 5 23
16 SPIE 1 -
17 De Gruyter 1 -
18 Wiley 1 1
19 Sage 1 -
20 Academic Pinnacle 1 1
Filtered sub-total 58 77
21 arXiv 11 23
22 Preprints - 1
Filtered total 69 101
Table 2: Distribution of scientific papers on Web of Science (WoS). (Filtered total is limited to the year range 2014 - 2024 and removes irrelevant results.)
# Publisher Multimodal Language
Unfiltered total 4 50
Corrected total 8 46
1 IEEE 4 5
2 Elsevier - 2
3 ACM 3 12
4 Springer - 5
5 NeurIPS - 1
6 Nature - 1
7 MDPI - 3
22 MLRP - 1
22 PKP 1 -
16 ACL - 10
13 Wiley - 2
17 AIAF - 1
12 Now - 1
Filtered total 8 44
Finally, for Scholar, we captured archived papers (e.g. arXiv) because, sometimes, their peer-reviewed versions exist but may not appear in the result. The summary statistics of the search are presented in Tables 1 and 2 while the references to the papers and the link to the search results are provided in the Appendix. The useful data from the reviewed papers (or primary studies) about datasets, AI models, and other contributions are then discussed in Sections 4.1, 4.2, and 5, respectively. We note that the more relevant papers turn up on the first page or at the top of the search while the quality of results degrade as one progresses through the pages or list. From the Tables, it can be observed that there are fewer pair-reviewed papers on multimodal models compared to language models.
4 Findings on Fairness and Bias in LMM s and LLM s
It is commonly agreed that AI models learn much of their bias from the data they are trained on and many datasets, especially those for pretraining, are from the Internet, which contains a diverse spectrum of content Wolfe and Caliskan (2022a). Tables 3 and 4 summarize some relevant datasets and the models beset by challenges of bias, respectively. All the 25 datasets identified have their challenges and by extension the 25 AI models which train on them. Some of these challenges include stereotypes, porn, misogyny, racial, gender, religious, cultural, age, and demographic biases.
@@ 4.1 LMM s
Table 3: Summary of Some Datasets and Their Fairness & Bias Challenges (Data in the lower part of the table are usually used in downstream tasks).
# Dataset Modality Some Challenges of Bias/Fairness
@@
10 The Pile Gao et al. (2020a) Text Religious, racial, gender biases Gao et al. (2020a)
@@
22 Multi30K Elliott et al. (2016) Text & Vision Racial bias Wang et al. (2022a)
23 MIMIC-IV Johnson et al. (2023) Text & Vision Ethnic, racial, marital status biases Meng et al. (2022a)
@@
25 Twitter corpus Huang et al. (2020b) Text Age, gender, racial biases Huang et al. (2020b)
Liang et al. (2021a) acknowledged that multimodal representations are challenging because they seek to integrate information from multiple areas in applications like robotics, finance, healthcare and more. However, multimodal AI has not enjoyed enough resources to study generalization across different modalities and the complexities of training. With regards to bias, Wolfe et al. (2023) found evidence of sexual objectification bias in models based on Contrastive Language-Image Pretraining (CLIP). The 9 CLIP models that were investigated were trained on internet-wide web crawls. CLIP is known to be quite accurate Radford et al. (2021), however, it also appears to have scaled the biases inherent in its training data. Also, Wolfe and Caliskan (2022a) found that more than Latino, Asian or Black, White persons are more associated with collective in-group words in embeddings from CLIP Radford et al. (2021), SLIP Mu et al. (2022), and BLIP Li et al. (2022), as measured with Embedding Association Tests (EATs). For a definitive assessment, their work would have benefited from additional experiments involving data of people outside the United States (US), since they used the Chicago Face Database (CFD)Ma et al. (2015). CFD is a dataset of 597 people recruited in the U.S. Similarly, Teo et al. (2024) found that Stable Diffusion exhibits gender bias on slight changes to its prompts.
Besides the work on text and text-visual multimodal systems or data, there are some work on audio-visual systems or data Fenu and Marras (2022). In the work by Fenu and Marras (2022), they perform comparative analysis on audio-visual speaker recognition systems, using fusion at the model step. They found that the highest accuracy and the lowest disparity across groups are achieved compared to unimodal systems.
In other works, Peña et al. (2023) evaluated AI-based recruitment for multimodal data but in a fictitious case study, which may limit its generalizability in real-world applications. Booth et al. (2021) performed a case study of automated video interviews and found that combining more than one modality increases bias and reduces fairness, similarly to what happens when scaling crawled data for training models. Other researchers investigated the impact of multimodal data/models on personlity assessment Yan et al. (2020), cyberbullying Alasadi et al. (2020), health records Meng et al. (2022a) and more Birhane et al. (2021); Zhang et al. (2022c); Ferrara (2023a); Cabello et al. (2023) The obvious challenges of bias in data has motivated some researchers for more attention in the ethics of data collection Weinberg (2022).
@@
# LMM s Modality Training Data Some Challenges of Bias/Fairness
1 VQGAN-CLIP Crowson et al. (2022) Text & Vision WIT & ImageNet Misogyny, racial bias Pagliai et al. (2024); Wolfe and Caliskan (2022a)
2 DALL-E 2 Ramesh et al. (2021) Text & Vision Conceptual Captions Occupational stereotypes, gender bias Ramesh et al. (2022); Mandal et al. (2023a)
@@
5 SLIP Mu et al. (2022) Text & Vision YFCC100M Racial bias Wolfe and Caliskan (2022a)
6 CLIP Radford et al. (2021) Text & Vision WIT Racial bias Radford et al. (2021); Wolfe and Caliskan (2022a)
7 BLIP Li et al. (2022) Text & Vision MS COCO, Conceptual Captions & LAION-400M Racial, gender & age biases Wolfe and Caliskan (2022a); Ruggeri and Nozza (2023)
8 PaliGemma Text & Vision WebLI, CC3M-35L , WIT Porn, offensive language333ai.google.dev/gemma/docs/paligemma/model-card
@@
11 BEiT Wang et al. (2023b) Text & Vision Conceptual 12M, ImageNet-21K, Wikipedia Gender, cultural biases Brinkmann et al. (2023)
@@
23 Aurora-M Nakamura et al. (2024) Text The Pile Offensive language, religious, racial, gender biases Gao et al. (2020a); Nakamura et al. (2024)
24 Mixtral-8x7B Jiang et al. (2024a) Text web text Stereotypes, racial, gender, occupational biases Jiang et al. (2024a); Aich et al. (2024)